40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 21 17:01:59 2022, Paco, Update, BHD, BH55 RFPD installed - part I IMG_3760.jpegIMG_3764.jpeg
    Reply  Thu Sep 22 19:51:58 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part II 
       Reply  Fri Sep 23 19:04:12 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part III 
          Reply  Thu Sep 29 18:01:14 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part IV 
             Reply  Fri Sep 30 18:30:12 2022, Anchal, Update, ASS, Model Changes 
                Reply  Wed Oct 12 14:01:23 2022, Radhika, Update, ASS, Model Changes YARM_ASS.png
                   Reply  Wed Oct 12 18:18:16 2022, rana, Update, ASS, Model Changes 
    Reply  Fri Sep 23 14:10:19 2022, Radhika, Update, BHD, BH55 RFPD installed - part I IMG_3766.jpegIMG_3770.jpegIMG_3773.jpeg
       Reply  Fri Sep 23 18:31:46 2022, rana, Update, BHD, BH55 RFPD installed - part I 
          Reply  Mon Sep 26 11:39:37 2022, Paco, Update, BHD, BH55 RFPD installed - part II PXL_20220926_175010061.MOTION.jpgBH55_Transimpedance_Measurement.pdf
             Reply  Thu Oct 6 11:12:14 2022, Anchal, Update, BHD, BH55 RFPD installation complete 
Message ID: 17155     Entry time: Fri Sep 23 14:10:19 2022     In reply to: 17150     Reply to this: 17156
Author: Radhika 
Type: Update 
Category: BHD 
Subject: BH55 RFPD installed - part I 

[Radhika, Paco, Anchal]

I placed a lens in the B-beam path to focus the beam spot onto the RFPD [Attachment 1]. To align the beam spot onto the RFPD, Anchal misaligned both ETMs and ITMY so that the AS and LO beams would not interfere, and the PD output would remain at some DC level (not fringing). The RFPD response was then maximized by scanning over pitch and yaw of the final mirror in the beam path (attached to the RFPD).

Later Anchal noticed that there was no RFPD output (C1:LSC-BH55_I_ERR, C1:LSC-BH55_Q_ERR). I took out the RFPD and opened it up, and the RF OUT SMA to PCB connection wire was broken [Attachment 2]. I re-soldered the wire and closed up the box [Attachment 3]. After placing the RFPD back, we noticed spikes in C1:LSC-BH55_I_ERR and C1:LSC-BH55_Q_ERR channels on ndscope. We suspect there is still a loose connection, so I will revisit the RFPD circuit on Monday. 

Attachment 1: IMG_3766.jpeg  348 kB  Uploaded Fri Sep 23 15:34:38 2022  | Hide | Hide all
IMG_3766.jpeg
Attachment 2: IMG_3770.jpeg  214 kB  Uploaded Fri Sep 23 17:42:50 2022  | Hide | Hide all
IMG_3770.jpeg
Attachment 3: IMG_3773.jpeg  209 kB  Uploaded Fri Sep 23 17:42:56 2022  | Hide | Hide all
IMG_3773.jpeg
ELOG V3.1.3-